### v1

# The Mollified (Discrete) Uniform Distribution and its Applications

Abstract
The mollified uniform distribution is rediscovered, which constitutes a “soft” version of the continuous uniform distribution. Important stochastic properties are derived and used to demonstrate potential fields of applications. For example, it constitutes a model covering platykurtic, mesokurtic and leptokurtic shapes. Its cumulative distribution function may also serve as the soft-clipping response function for defining generalized linear models with approximately linear dependence. Furthermore, it might be considered for teaching, as an appealing example for the convolution of random variables. Finally, a discrete type of mollified uniform distribution is briefly discussed as well.
Key words: convolution; discrete uniform distribution; kurtosis; mollified uniform distribution; soft-clipping regression.
1 Introduction
This article deals with a surprisingly little-known distribution, namely the mollified uniform distribution, the cumulative distribution function (cdf) of which is also known as the soft clipping function. It has been occasionally used in practice, e. g., as an output or activation function for artificial neural networks (ANNs), see Klimek & Perelstein (2020), or as a response function for time series models (Weiß & Jahn, 2022), and it has been discussed in forums such as “Cross Validated” (https://stats.stackexchange.com/q/332558, accessed on 16 February 2024). But there is hardly any literature about the distribution itself and its potential applications. The low level of awareness is also surprising in view of the fact that the mollified uniform distribution would be well suited for teaching, namely as a non-trivial (and quite appealing) example for the convolution of random variables. Accordingly, this article intends to dispel the aforementioned shortcomings and to present possible applications of the mollified uniform distribution.
The idea of mollifying the (standard) uniform distribution $\textup{U}(0,1)$ dates back to Friedrichs (1944), who proposed “a certain class of smoothing operators approximating unity” (p. 132), namely by adding an independent and centered random variable $L$ (the mollifier) with very low dispersion to the uniform random variable $U$ (for properties of the uniform distribution, see Johnson et al. (1995, Ch. 26)). While Friedrichs (1944) originally focused on mollifiers $L$ with compact support, it is more convenient to consider random variables $L$ on the full set $\mathbb{R}$ of reals, having a symmetric distribution around zero and arbitrarily low dispersion. These could be a normally distributed $L\sim\textup{N}(0,\sigma)$ with (low) standard deviation $\sigma>0$ (see Johnson et al., 1994, Ch. 13), or a Laplace distributed $L\sim\textup{La}(0,b)$ with scale parameter $b>0$ (see Johnson et al., 1995, Ch. 24). But the most suitable mollifier (e. g., in view of simplicity of the resulting cdf) appears to be the logistic one, namely $L\sim\textup{L}(0,c)$ with scale parameter $c>0$ (see Johnson et al., 1995, Ch. 23). Independent of the specific choice of $L$ (although we often consider $L\sim\textup{L}(0,c)$ in the sequel), the mollified uniform distribution is defined to be the distribution of $X=U+L$, i. e., its probability density function (pdf) is obtained as the convolution of the mollifier’s pdf $f_{L}(\cdot)$ and the uniform’s pdf $f_{U}(\cdot)$. Here, the latter equals $f_{U}(x)=\mathds{1}_{(0;1)}(x)$, where the indicator function $\mathds{1}_{A}(x)$ equals 1 (0) if $x\in A$ ($x\not\in A$). Hence, the pdf $f(\cdot)$ of $X$ just equals
$f(x)\ =\ \int_{-\infty}^{\infty}f_{L}(x-y)\,f_{U}(y)\,\textup{d}y\ =\ \int_{0}^{1}f_{L}(x-y)\,\textup{d}y\ =\ F_{L}(x)-F_{L}(x-1),$ (1.1)
where $F_{L}(\cdot)$ denotes the mollifier’s cdf. Equation (1.1) gets particularly simple if $L\sim\textup{L}(0,c)$ where $F_{L}(x)=\big{(}1+\exp(-x/c)\big{)}^{-1}$, namely
$f(x)\ =\ \frac{1}{1+\exp(-\frac{x}{c})}\,\frac{1-\exp(-\frac{1}{c})}{1+\exp(\frac{x-1}{c})},$ (1.2)
which is also plotted in Figure 1 (a). Note that the cdf in (b) and the quantile function (qf) in (c) are introduced later in Section 4.
[…]
Figure 1: Logistic-mollified uniform distribution with scale parameter $c$: plots of (a) pdf, (b) cdf, and (c) qf for different $c$.
Looking at the “rounded edges” for small $c$ in Figure 1 (a), it gets clear that also the denomination as a “soft uniform distribution” would be appropriate. This term, however, has been used so far for a different construction (see Yang & Rannala, 2006): the soft-uniform’s pdf is defined in a piecewise manner (continuous, but not differentiable), where the rectangular part is expanded by a power decay on the left side and an exponential decay on the right side. At this point, it is worth mentioning that also a (centered) asymmetric distribution could be used for mollification according to (1.1), although the aforementioned symmetric mollifiers are probably more convenient for practice.
In what follows, we briefly discuss some possible application scenarios for the mollified uniform distribution. In Section 2, we study the mollified uniform’s moment properties and show that it may serve as a model that comprises a wide range of kurtosis levels, from platykurtic over mesokurtic up to leptokurtic shapes (i. e., kurtosis $<3$ or $=3$ or $>3$, respectively). It should also be pointed out that the required derivations are elementary, just making use of the convolution structure (1.1), such that they may be considered as a neat example for teaching. Section 3, in turn, constitutes a brief digression to the discrete case: a discrete mollification of the discrete uniform distribution is proposed, which would again be suitable for classroom presentation. Section 4 then moves on to the mollified uniform’s cdf and qf (also recall Figure 1), and explains how the distribution might be used for expressing “nearly linear” generalized linear models (GLMs), e. g., for binary, ordinal, or bounded quantitative data. Finally, Section 5 concludes and outlines issues for future research.
2 Moments and Kurtosis
The raw moments of the uniform $U$ are easily calculated as $E[U^{n}]=\frac{1}{n+1}$ for $n\in\mathbb{N}_{0}=\{0,1,\ldots\}$. Since a mollified uniform random variable $X$ equals the convolution $X=U+L$, recall (1.1), its moments follow from the moments of the mollifier $L$ by applying the binomial sum formula:
$E[X^{n}]\ =\ \sum_{k=0}^{n}\binom{n}{k}\,E[U^{n-k}]\,E[L^{k}]\ =\ \sum_{k=0}^{n}\binom{n}{k}\,\frac{1}{n-k+1}\,E[L^{k}].$ (2.1)
If using a symmetric (centered) mollifier, moments $E[L^{k}]$ of odd order $k$ are equal to zero. Moment expressions for even order $k$, in turn, can be found in Johnson et al. (1994, 1995). For example, $E[L^{k}]=b^{k}\,k!$ if $L\sim\textup{La}(0,b)$, or $E[L^{k}]=\sigma^{k}\,(k-1)!!$! if $L\sim\textup{N}(0,\sigma)$, where the double factorial of the odd number $k-1$ is the product of all odd numbers between 1 to $k-1$. The even-order moments of the logistic mollifier $L\sim\textup{L}(0,c)$ can be generally expressed by using so-called Bernoulli numbers; in particular, $E[L^{2}]=\frac{\pi^{2}\,c^{2}}{3}$ and $E[L^{4}]=\frac{7\pi^{4}\,c^{4}}{15}$.
In the case of a symmetric mollifier $L$, it follows that $X$ is symmetric around its mean $E[X]=\frac{1}{2}$. So it is more appropriate to look at central moments of $X$ rather than at its raw moments, where obviously $E\big{[}(X-\frac{1}{2})^{n}\big{]}=0$ if $n$ is odd. For even $n$, the central moments follow from (2.1) by again applying the binomial sum:
$E\big{[}(X-\tfrac{1}{2})^{n}\big{]}\ =\ \sum_{k=0}^{n}\binom{n}{k}\,\frac{E[X^{k}]}{(-2)^{n-k}}.$ (2.2)
As an example, for $n=2$, it follows together with (2.1) that $E\big{[}(X-\tfrac{1}{2})^{2}\big{]}=\frac{1}{12}+E[L^{2}]$, i. e., the variance satisfies $V[X]=\frac{1}{12}+V[L]$, which is also implied by Bienaymé’s identity.
We shall apply formulae (2.2) and (2.1) to compute the kurtosis of $X$, i. e., $\kappa(X)=E\big{[}(X-\tfrac{1}{2})^{4}\big{]}/E\big{[}(X-\tfrac{1}{2})^{2}\big{]}{}^{2}$. For a comprehensive discussion of kurtosis and possible (mis)interpretations, see Westfall (2014). A simple calculation leads to
$E\big{[}(X-\tfrac{1}{2})^{4}\big{]}\ =\ \frac{1}{80}+\frac{1}{2}\,E[L^{2}]+E[L^{4}]\ =\ \frac{1}{80}+\frac{1}{2}\,V[L]+V[L]^{2}\cdot\kappa(L),$
where $\kappa(L)=E[L^{4}]/E[L^{2}]^{2}$ is the mollifier’s kurtosis (recall that $E[L]=0$). For the example cases discussed before, one gets
$\begin{array}[]{lll}V[L]\ =\ \sigma^{2}&\text{and}\quad\kappa(L)\ =\ 3&\text{if }L\sim\textup{N}(0,\sigma),\\[4.30554pt] V[L]\ =\ \frac{\pi^{2}}{3}\,c^{2}&\text{and}\quad\kappa(L)\ =\ \frac{21}{5}&\text{if }L\sim\textup{L}(0,c),\\[4.30554pt] V[L]\ =\ 2b^{2}&\text{and}\quad\kappa(L)\ =\ 6&\text{if }L\sim\textup{La}(0,b).\end{array}$ (2.3)
So finally, assuming a constant $\kappa(L)$ like in (2.3),
$\kappa(X)\ =\ \frac{\frac{1}{80}+\frac{1}{2}\,V[L]+V[L]^{2}\cdot\kappa(L)}{\big{(}\frac{1}{12}+V[L]\big{)}^{2}}\quad\to\ \left\{\begin{array}[]{ll}\frac{144}{80}=\frac{9}{5}&\text{if }V[L]\to 0,\\[4.30554pt] \kappa(L)&\text{if }V[L]\to\infty.\end{array}\right.$ (2.4)
Hence, modifying the mollifier’s variance, we can move between the uniform’s kurtosis $\frac{9}{5}<3$ (platykurtic) and the mollifier’s kurtosis $\kappa(L)$. In Figure 2 (a), the kurtosis $\kappa(X)$ is plotted against the mollifiers’ variance $V[L]$ for the cases (2.3). The dotted lines indicate the kurtosis of $\textup{U}(0,1)$, $\textup{L}(0,c)$, and $\textup{La}(0,b)$, respectively, whereas the solid line refers to the normal’s kurtosis 3. In particular, the solid line expresses the mesokurtic case, while kurtosis values below (above) the solid line correspond to the platykurtic (leptokurtic) case. If using a normal mollifier, the mollified uniform is always platykurtic, whereas for logistic or Laplace mollifiers, one can move across all three cases. Furthermore, recalling the pdf in Figure 1 (a), the mollified uniform distribution might also serve as an illustrative example for the main message in Westfall (2014), namely that “kurtosis tells you very little about the peak or center of a distribution” (p. 194).
[…]
Figure 2: Kurtosis of mollified uniform against mollifier’s variance in (a). Skellam-mollified discrete uniform’s pmf with $m=10$ plotted for (b) $\lambda=0.1$ and (c) $\lambda=0.5$.
3 Intermezzo: Mollified Discrete Uniform Distribution
Before continuing with the mollified uniform’s cdf and corresponding applications in Section 4, let us have a brief look at the discrete case. In complete analogy to (1.1), one can define the mollified discrete uniform distribution as the convolution $Y=V+Z$, where $V$ follows the (standard) discrete uniform distribution $\mathcal{U}\{1,m\}$ with $m\in\mathbb{N}=\{1,2,\ldots\}$ and $m\geq 2$, and where $Z$ is symmetrically distributed around 0 with support =$\mathbb{Z}=\{\ldots,-1,0,1,\ldots\}$. Here, $P(V=k)=\frac{1}{m}$ for $k\in\{1,\ldots,m\}$ and 0 otherwise. The probably most well-known candidate for $Z$ is the Skellam distribution (see Johnson et al., 2005, Section 4.12.3), which is itself a convolution of two random variables. More precisely, $Z\sim\textup{Sk}(\lambda_{1},\lambda_{2})$ can be generated as the difference $Z=Y_{1}-Y_{2}$ of the two independent Poisson variates $Y_{1}\sim\textup{Poi}(\lambda_{1})$ and $Y_{2}\sim\textup{Poi}(\lambda_{2})$. Its mean and variance are given by $\lambda_{1}-\lambda_{2}\in\mathbb{R}$ and $\lambda_{1}+\lambda_{2}>0$, respectively. Thus, to be suitable as a mollifier, we consider the centered symmetric case $Z\sim\textup{Sk}(\lambda,\lambda)$ such that $V[Z]=2\lambda$ and $E[Z^{k}]=0$ for odd $k$.
The probability mass function (pmf) of $Y=V+Z$ is computed in analogy to (1.1), namely as
$\begin{array}[]{@{}rl}P(Y=y)\ =&\sum_{k=1}^{m}P(Z=y-k)\ P(V=k)\ =\ \tfrac{1}{m}\,\sum_{k=1}^{m}P(Z=y-k)\\[4.30554pt] =&\tfrac{1}{m}\,\big{(}P(Z\leq y-1)-P(Z\leq y-m-1)\big{)}\qquad\text{for }y\in\mathbb{Z}.\end{array}$ (3.1)
Illustrative plots of the pmf if using a Skellam mollifier are provided by parts (b) and (c) of Figure 2. Moment calculations for the mollified discrete uniform are done in nearly the same way as in Section 2, as we are still concerned with a convolution of the form $Y=V+Z$. One just has to adapt the uniform’s moments in (2.1), namely to $E[V^{n}]=\frac{1}{m}\,\sum_{k=1}^{m}k^{n}$, which can be generally computed by using Faulhaber’s formula. In particular, one gets the following (central) moments of $V$:
$E[V]\ =\ \tfrac{1}{2}\,(m+1),\quad E\big{[}(V-\tfrac{m+1}{2})^{2}\big{]}\ =\ \tfrac{1}{12}\,(m^{2}-1),\quad E\big{[}(V-\tfrac{m+1}{2})^{4}\big{]}\ =\ \tfrac{1}{240}\,(3m^{4}-10m^{2}+7).$
The (symmetrically) mollified discrete uniform distribution has mean $\frac{m+1}{2}$ as well, the odd central moments equal zero, whereas the even central moments follow from
$E\big{[}(Y-\tfrac{m+1}{2})^{n}\big{]}\ =\ \sum_{k=0}^{n}\binom{n}{k}\,E\big{[}(V-\tfrac{m+1}{2})^{n-k}\big{]}\,E[Z^{k}].$ (3.2)
Here, we have non-zero summands only for even $k$. For the kurtosis, we end up with
$\kappa(Y)\ =\ \frac{\frac{1}{240}\,(3m^{4}-10m^{2}+7)+\frac{1}{2}\,(m^{2}-1)\,V[Z]+V[Z]^{2}\cdot\kappa(Z)}{\big{(}\frac{1}{12}\,(m^{2}-1)+V[Z]\big{)}^{2}}.$ (3.3)
If using the leptokurtic Skellam mollifier having $\kappa(Z)=3+\frac{1}{2\lambda}$, then $\kappa(Y)$ tends to the platykurtic value $\kappa(V)=\frac{3}{5}\,(3-\frac{4}{m^{2}-1})$ for $V[Z]\to 0$. But the mollified discrete uniform distribution may also become mesokurtic, namely if $\lambda=\frac{1}{240}\,(m^{4}-1)$, or even slightly leptokurtic: the maximal kurtosis is attained for $\lambda=\frac{1}{120}\,(m^{4}+5m^{2}-6)$ and takes the value $33+\frac{30}{(m^{2}-1)(m^{2}+11)}>$ 3. Here, the latter result can be derived by simple calculus.
4 Soft-clipping Regression Models
Let us return to the continuous mollified uniform distribution. The cdf $F(x)$ of the mollified uniform distribution is generally obtained by integrating $f(x)=F_{L}(x)-F_{L}(x-1)$ according to (1.1). For the Laplace mollifier, however, this leads to a piecewise expression, whereas the cdf for the normal mollifier requires non-elementary functions. More precisely, let $\varphi(x)$ and $\Phi(x)$ denote the pdf and cdf of $\textup{N}(0,1)$. Then, using that $\varphi^{\prime}(x)=-x\cdot\varphi(x)$, one can show that an antiderivative of $\Phi(x)$ is given by $\varphi(x)+x\cdot\Phi(x)$, and this allows to express the cdf in case of the normal mollifier. By contrast, quite simple expressions for cdf $F$ and qf $F^{-1}$ are obtained if using the logistic mollifier. Then,
$F(x)\ =$ $\displaystyle c\,\ln\left(\frac{\exp(\frac{x}{c})+1}{\exp(\frac{x-1}{c})+1}\right)$ (4.1)
[…]
The second expression (see Weiß & Jahn, 2022) allows for a numerically stable implementation by using the log1p function, which is offered by common programming languages. In particular, it shows that $F(x)$ approaches the piecewise linear function $\min\{1,\max\{0,x\}\}$ for $c\to 0$, also see Figure 1 (b), which is the cdf of the uniform distribution. Here, $\min\{1,\max\{0,x\}\}$ is also known as the clipped rectified linear unit (cReLU) function (see Cai et al., 2017), so $F(x)$ constitutes a smooth version of it and is, thus, referred to as the soft clipping function $\textup{sc}_{c}(x)$ by Klimek & Perelstein (2020). The mollified uniform’s qf $F^{-1}:(0;1)\to\mathbb{R}$, in turn, equals
$F^{-1}(q)\ =\ c\,\ln\biggl{(}\frac{\exp(\frac{q}{c})-1}{1-\exp(\frac{q-1}{c})}\biggl{)}\ =\ q\,+\,c\,\ln\!\big{(}1-\exp(-\tfrac{q}{c})\big{)}\,-\,c\,\ln\!\big{(}1-\exp(-\tfrac{1-q}{c})\big{)},$ (4.2)
see Weiß & Jahn (2022), where the second expression is again suitable for the log1p function. A plot of the qf is provided by Figure 1 (c).
Besides probabilistic computations, the cdf $\textup{sc}_{c}(\cdot)$ from (4.1) and qf $\textup{sc}_{c}^{-1}(\cdot)$ from (4.2) can be used as response and link function, respectively, for regression analysis, namely for defining soft-clipping GLMs, also see Klimek & Perelstein (2020); Weiß & Jahn (2022). More precisely, soft-clipping GLMs are relevant for modeling binary, ordinal, or bounded quantitative response variables (e. g., binomial counts or beta-distributed outcomes). In all these cases, the logit GLM (i. e., with $\textup{L}(0,1)$’s qf as the link function) and probit GLM (using $\textup{N}(0,1)$’s qf) are most common yet, see McCullagh & Nelder (1989), but the soft-clipping GLM could constitute an interesting alternative. For illustration, consider the simple scenario where the response variable $Y$ is conditionally binomially distributed depending on some explanatory variable $x$, say $Y|x\sim\textup{Bin}(n,\pi_{x})$. In a GLM setup with link function $g:(0;1)\to\mathbb{R}$ and response function $h=g^{-1}$, one assumes that $g(\pi_{x})=a+b\,x$ and $\pi_{x}=h(a+b\,x)$, respectively. If using the logit GLM with $h(x)=\big{(}1+\exp(-x)\big{)}^{-1}$, it follows that the log-odds is linear in $x$, namely $\ln\!\big{(}\pi_{x}/(1-\pi_{x})\big{)}=a+b\,x$, but the relation between $Y$ and $x$ is highly non-linear. The latter also holds true for probit modeling, i. e., if $h=\Phi$.
If the relation between $Y$ and $x$ is close to linearity, another link is necessary. Using the identity link (i. e., assuming the exact linearity $\pi_{x}=a+b\,x$) is often not feasible in practice as strict constraints on the parameters $a,b$ are necessary to ensure that $a+b\,x\in(0;1)$ always holds. Parameter constraints are dispensable if using the cReLU response instead, but this choice (although often used for ANNs) has two disadvantages: cReLu is not differentiable such that common asymptotics of maximum likelihood estimation do not hold, and $\pi_{x}$ might become equal to 0 or 1 which leads to a degenerate binomial distribution. These problems are solved if using a soft-clipping response function $\textup{sc}_{c}(\cdot)$ instead, as this is infinitely differentiable and attains values only in the open interval $(0;1)$. In particular, (4.1) is close to (piecewise) linearity provided that $c$ is chosen sufficiently small, e. g., $c=0.01$ in view of Figure 1 (b).
Example 1
Let us illustrate the above discussion by a small simulation experiment. First, samples of 100 explanatory variables $x$ are generated according to $\textup{Bin}(n,0.5)$ with $n=30$. Then, $\pi_{x}=\textup{sc}_{c}(a+b\,\frac{x}{n})$ is computed for each $x$, with $a=0.4$ and $b=-0.6$, and the response $Y$ is generated from $\textup{Bin}(n,\pi_{x})$. Finally, an ordinary linear model is fit to the sample of pairs $(Y,x)$, and all steps were repeated for 1,000 times.
Analyzing these 1,000 replicates, it gets clear that soft clipping is indeed necessary for computing $\pi_{x}$, because if we would use $\pi_{x}=a+b\,\frac{x}{n}$, then in 995 simulation runs, at least one $\pi_{x}$ would not be positive (in the mean, $a+b\,\frac{x}{n}\leq 0$ happens about five times per sample). Next, let us look at the estimates for the fitted linear model $E[Y]=\alpha+\beta\,x$. The sample mean across the $\hat{\alpha}$ equals $\approx 11.7$, the one across $\hat{\beta}$ equals $\approx-0.58$. Under the exact linearity $\pi_{x}=a+b\,\frac{x}{n}$, we would expect that $E[Y|x]=n(a+b\,\frac{x}{n})=na+b\,x$, i. e., $\alpha=12$ and $\beta=-0.6$. Looking at the aforementioned means of estimates, we recognize a close agreement, i. e., the soft-clipping GLM that was used for data generation indeed behaves very similar to a truly linear model, but avoids the problem of inadmissible $\pi_{x}$-values.
5 Conclusions
In this article, the (continuous) mollified uniform distribution was rediscovered, and a novel discrete counterpart was proposed as well. Both types of mollified uniform distribution are defined as the convolution of the (discrete) uniform distribution and a centered symmetric mollifier. Relevant stochastic properties are computed by elementary calculations such that the mollified uniform distributions might serve as neat examples in teaching. The mollified uniform distribution also constitutes a model that may cover a wide range of kurtosis levels. Furthermore, its cdf and qf, respectively, are used for the defining soft-clipping regression models, which can be adapted to binary, ordinal, and bounded quantitative outcomes. Here, a couple of future research directions become clear. Up to now, the soft-clipping GLM approach was only investigated in rather special setups, namely as the output function of an ANN in order to generate bounded real-valued outcomes in the unit interval $(0;1)$, see Klimek & Perelstein (2020), or for modeling time series of bounded counts (see Weiß & Jahn, 2022). But applications to, e. g., ordinary count regression (without the time aspect), to binary data, or ordinal data (as cumulative GLMs) appear to not have been considered so far.